OpenStudy (anonymous):

What are the coordinates of the center of a circle whose diameter has the endpoints (1, 4) and (-3, -3)? (-2, 1) (-1, 0.5) (-1, 0) (2, 3.5)

OpenStudy (logan13):

@Nicole28prow

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

the center is at the midpoint of the diameter, so you can use the midpoint formula \[\left( \frac{ x_1+x_2 }{ 2 },\frac{ y_1+y_2 }{ 2 } \right)\] Use your point for the x's and y's